OpenStudy (alex6799):

Evaluate the expression for the given value of the variable(s). -x 2 - 4x; x = -3 A. 11 B. 3 C. -17 D. -1

OpenStudy (alex6799):

\[-x ^{2}-4x; x=-3\]

OpenStudy (alex6799):

\[-3^{2}-4*-3\]

OpenStudy (alex6799):

i keep getting 21 but 21 isn't an answer choice

OpenStudy (solomonzelman):

For example, if I want to evaluate: \(\color{#000000}{ \displaystyle -x^2-4x;\quad {\rm at}~~~x=-2 }\) Plug in x=-2 into the expression. \(\color{#000000}{ \displaystyle -(-2)^2-4(-2)= -(4)-(-8) =-4+8=-4 }\)

OpenStudy (alex6799):

\[-3^{2}-4*-3\] \[(-3^2)-4(-3)\] \[(9)-4(-3)\] right so far?

OpenStudy (solomonzelman):

No, you are missing something

OpenStudy (solomonzelman):

I will point out the correction in red. \(\color{#000000}{ \displaystyle \color{red}{-(-3)^2 }-4(-3) }\)

OpenStudy (alex6799):

ooooooh so -9

OpenStudy (alex6799):

so the answer is b?

OpenStudy (solomonzelman):

Yes, B is the correct answer.

OpenStudy (alex6799):

thank you!

OpenStudy (solomonzelman):

\(\color{#0a00c0}{ \displaystyle -(-3)^2-4(-3) =-(9)-(-12) =-9+12=3 }\)